Operations Report – February 16th

Crew IKEA Operations Report 16 Feb 2018

SOL: 6

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: Nothing new to report

Generator (hours run): Off at 8:30 am, on at 6 pm. Turns out the

power system works just fine with lots of welding happening!

Solar— 98% SOC

Diesel – 72 %

Propane – 46 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 10 gallons.

Water (trailer) – 0 gallons.

Water (static) – 400 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: no

Oil Added? n/a

ATV Fuel Used: n/a Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

Deimos rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 96

Currently charging: Yes

Sojourner rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: Yes

Hours: 20.2

Beginning charge: 100

Ending charge: 75

Currently charging: No

Opportunity rover used: Yes

Hours: 14.9

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: No

HabCar used and why, where? Yes. To town for water for the

aquaponics, packages, searching for the trailer hitch in storage. To

Valley of the Stars for fieldwork. To town to pack food. To town to

get gas for the ATVs and more water.

General notes and comments: How did it get to be Friday already?

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Installations of the stairs is almost

complete! We will have photos tomorrow.

Summary of GreenHab operations: Hydroponics (no fish) set up complete

and water added. We will get it circulating tomorrow. We also

realized that we have two kinds of tomatoes, regular sized and cherry.

Atila and I each had two of the best cherry tomatoes I have ever

eaten.

Summary of ScienceDome operations: We did some lab work this morning.

Summary of RAM operations: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing at this time